A ladder is resting against a wall. The top of the ladder touches the wall at a height of 9ft. Find the length of the ladder if the length is 3ft more than its distance from the wall.
Answer by mananth(16946)

wall ladder distance =x
length of ladder = x+3
height touching the wall = 9
(x+3)^2=x^2+9^2
x^2+6x+9=x^2+81
6x=81-9
6x=72
/6
x=12 feet
ladder ===>x+3= 12+3=15 feet
m.ananth@hotmail.ca